OpenStudy (anonymous):

You and some friends have started your own company. After the first few months, the profits are rolling in. It is time to start thinking about putting your money to work for you. You decide that investing $5,000 into some Certificates of Deposit (CDs) would be a beneficial move. With a CD, you lend your money to a third party, and after a set time, your money is paid back with interest. Before you start investing the company's money this way, you need to pitch it to your friends.

OpenStudy (anonymous):

1. Research the highest interest rate (APY—annual percent yield) for 2-year and 5-year CDs. Document the company's name, interest rate, and minimum investment. The minimum investment must be less than or equal to $5,000. f(x) = 5000(1.2 for example)x You decide that investing $5,000 into some Certificates of Deposit (CDs) would be a beneficial move. Now u would just plug in the years as x...for example: f(1) = 5000 x 1.2 f(1) = 6000 f(2) = 5000(1.2)2 F(2) = 5000 x 2.4 f(2) = 12000 (your investing $5000) f(1) = 11500 f(2) = 5000(2.3)2 F(2) = 5000 x 4.6 f(2) = 23000 (your investing $5000) 2. Create the functions that represent the 2-year and 5-year CDs with your $5,000 investment. Use these functions to determine the amount you will be paid when the CD matures (the length of time for the specific CD). Show your work. f(2) = 12000 3. An investor comes to your office. He says that if you give him the $5,000, he will add on an additional $50 each year to what he owes you. Create the function for this investor's plan. f(t) = 5000 + 50t 4. Create a table showing the value of the two CDs and the investor's plan for 5 years. Year 1 Year 2 Year 3 Year 4 Year 5 2-year CD 6050 5-year CD Investor

OpenStudy (anonymous):

@eHoaX can you help with 4? 1, 2, and 3 are already done

OpenStudy (anonymous):

I'm thinking...

OpenStudy (anonymous):

@Peabody ok

OpenStudy (anonymous):

wait, no one has answered you yet!?

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

the 6050 is part of the answer for number 4

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@Peabody i tried asking hero yesterday, and he's probably busy

OpenStudy (anonymous):

sorry I have to be honest I don't kow..

OpenStudy (anonymous):

repost thequestion and someone will probably help

OpenStudy (anonymous):

@Peabody ok, this is my third time posting this question

OpenStudy (anonymous):

wow! :(

OpenStudy (anonymous):

@Boblovesmath Can you help with number 4? 1, 2, and 3 are already answered

OpenStudy (anonymous):

the 6050 is part of the answer

OpenStudy (anonymous):

from 1 you have two functions. The first one is for the 2 year CD and the second one is for the 5-year CD. x -value is the year and the output is the benefits (7-value)

OpenStudy (anonymous):

@Boblovesmath i didn't understand that, can you explain it a little simpler?

OpenStudy (anonymous):

You have two cases here you have the 2 year issue and the 5-year. Gather all of the information you have from the previous parts of the question and let x-value be the number of years and y-value by the output. Note that here you should have two tables one for the 2 year and the other for the 5-year. I can't help you in this case because I don't understand the steps that you used to get the solution. Hope this helped mate.

OpenStudy (anonymous):

@Boblovesmath thanks

OpenStudy (anonymous):

Glad to help :)